# Restricted $L_{\infty}$-algebras and a derived Milnor-Moore theorem

### v1

@@ Abstract.
For every stable presentably symmetric monoidal $\infty$-category ${\mathcal{C}}$ we use the Koszul duality between the spectral Lie operad and the cocommutative cooperad to construct an enveloping Hopf algebra functor ${\mathcal{U}}:\mathrm{Alg}_{\mathrm{Lie}}({\mathcal{C}})\to{\mathrm{Hopf}}({\mathcal{C}})$ from spectral Lie algebras in ${\mathcal{C}}$ to cocommutative Hopf algebras in ${\mathcal{C}}$ left adjoint to a functor of derived primitive elements. We prove that if ${\mathcal{C}}$ is a rational stable presentably symmetric monoidal $\infty$-category, the enveloping Hopf algebra functor ${\mathcal{U}}:\mathrm{Alg}_{\mathrm{Lie}}({\mathcal{C}})\to{\mathrm{Hopf}}({\mathcal{C}})$ is fully faithful. We conclude that Lie algebras in ${\mathcal{C}}$ are algebras over the monad underlying the adjunction ${\mathrm{T}}\simeq{\mathcal{U}}\circ{\mathrm{Lie}}:{\mathcal{C}}\rightleftarrows\mathrm{Alg}_{\mathrm{Lie}}({\mathcal{C}})\rightleftarrows{\mathrm{Hopf}}({\mathcal{C}}),$ where ${\mathrm{Lie}}$ is the free Lie algebra and ${\mathrm{T}}$ is the tensor algebra. For general ${\mathcal{C}}$ we introduce the notion of restricted $L_{\infty}$-algebra as an algebra over the latter adjunction. For any field ${\mathrm{K}}$ we construct a forgetful functor from restricted Lie algebras in connective ${\mathrm{H}}({\mathrm{K}})$-modules to the $\infty$-category underlying a right induced model structure on simplicial restricted Lie algebras over ${\mathrm{K}}$.
@@ 1. Introduction
In the seminal paper [24] Milnor and Moore prove that every graded Lie algebra over a field of characteristic zero arises as the primitive elements of its universal enveloping Hopf algebra [24, Proposition 5.17.]. From this result Milnor and Moore reveal a fundamental connection between Lie algebras and Hopf algebras over any field $K$ of characteristic zero: forming the universal enveloping Hopf algebra and forming the primitive elements are part of an adjunction
(1) ${\mathcal{U}}:{\mathrm{Lie}}_{K}\rightleftarrows{\mathrm{Hopf}}_{K}:{\mathrm{Prim}}$
between the categories of graded Lie algebras and graded Hopf algebras over $K$. By the Milnor-Moore theorem the unit of this adjunction is invertible, which for formal reasons implies the following theorem:
Theorem 1.1.
[24, Theorem 5.18.] Let $K$ be a field of characteristic zero. The unit of adjunction (1) is an isomorphism.
Moreover Milnor-Moore [24, Theorem 5.18.] characterize the Hopf algebras in the essential image of the enveloping Hopf algebra ${\mathcal{U}}$ as the primitively generated Hopf algebras, those graded Hopf algebras whose underlying graded algebra is generated by the subspace of primitive elements.
@@
(2) ${\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\mathsf{Ho}({\mathcal{S}}^{\geq 2}_{{\mathbb{Q}}})}$$\scriptstyle{{\mathcal{C}}}$$\scriptstyle{{\mathcal{L}}}$${\mathsf{Ho}({\mathrm{coAlg}}({\mathbb{Q}})^{\geq 2})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathsf{Ho}({\mathrm{Lie}}({\mathbb{Q}})^{\geq 1})}$
@@
Quillen builds the Lie model of a simply connected space $X$ in three steps. First he forms Kan’s loop group, a simplicial group that endows the loop space $\Omega({\mathrm{X}})$ with a group structure. As a second step he applies rational chains to Kan’s loop group to obtain a simplicial rational cocommutative Hopf algebra. Finally, he takes the primitive elements of this simplicial Hopf algebra to obtain a simplicial rational Lie algebra, which corresponds to a rational differentially graded Lie algebra via Dold-Kan’s correspondence. Quillen’s construction guarantees that the homology of the Lie model identifies with the primitive elements of the rational homology of the loop space, in symbols $H_{*}({\mathcal{L}}(X))\cong{\mathrm{Prim}}(H_{*}(\Omega(X);{\mathbb{Q}})).$ Since $X$ is simply connected, $H_{*}(\Omega(X);{\mathbb{Q}})$ is connected, which ensures that the graded Hopf algebra $H_{*}(\Omega(X);{\mathbb{Q}})$ is primitively generated. As a consequence the Milnor-Moore theorem identifies the enveloping Hopf algebra of $\pi_{*}(\Omega(X))\otimes{\mathbb{Q}})\cong H_{*}({\mathcal{L}}(X))\cong{\mathrm{Prim}}(H_{*}(\Omega(X);{\mathbb{Q}}))$ with $H_{*}(\Omega(X);{\mathbb{Q}})$ and so reveals a fundamental relationship between rational homotopy and rational homology of a loop space.
From modern perspective Quillen’s equivalence (2) between the Lie model and the coalgebra model is an instance of Koszul duality. Koszul duality between Lie algebras and cocommutative coalgebras or rather the homotopically refined notions of $L_{\infty}$-algebras and ${\mathbb{E}}_{\infty}$-coalgebras arises from the operadic Koszul duality between the shifted Lie operad ${\mathrm{Lie}}(1)$ and the non-unital cocommutative cooperad [9, Theorem 6.8.]. It takes the form of an adjunction
(3) $\mathrm{Alg}_{{\mathrm{Lie}}(1)}({\mathbb{Q}})\rightleftarrows{\mathrm{coAlg}}_{{\mathbb{E}}_{\infty}}({\mathbb{Q}}):{\mathrm{Prim}}$
between rational ${\mathrm{Lie}}(1)$-algebras and rational coaugmented ${\mathbb{E}}_{\infty}$-coalgebras, where the right adjoint assigns a derived version of primitive elements and the left adjoint takes the shifted Chevalley-Eilenberg complex computing Lie homology. By a theorem of Ching-Harper [4, Theorem 1.2.] adjunction (3) restricts to an equivalence between simply connected objects. Since a ${\mathrm{Lie}}(1)$-algebra structure on a rational chain complex is a $L_{\infty}$-structure on the negative shift, adjunction (3) restricts to an equivalence.
(4) $\mathrm{Alg}_{{\mathrm{Lie}}}({\mathbb{Q}})^{\geq 1}\simeq{\mathrm{coAlg}}_{{\mathbb{E}}_{\infty}}({\mathbb{Q}})^{\geq 2}$
between connected rational $L_{\infty}$-algebras and simply connected rational coaugmented ${\mathbb{E}}_{\infty}$-coalgebras, recovering Quillen’s result via the machinery of Koszul duality.
The interpretation of equivalence (2) as Koszul duality suggests a construction of the derived universal enveloping Hopf algebra purely expressed in terms of Koszul duality: derived Hopf algebras are group structures on ${\mathbb{E}}_{\infty}$-coalgebras and adjunction (3) descends to an adjunction on group objects
(5) ${\mathrm{Grp}}(\mathrm{Alg}_{{\mathrm{Lie}}(1)}({\mathbb{Q}}))\rightleftarrows{\mathrm{Hopf}}({\mathbb{Q}}):={\mathrm{Grp}}({\mathrm{coAlg}}_{{\mathbb{E}}_{\infty}}({\mathbb{Q}})):{\mathrm{Prim}}.$
@@
By Corollary 5.8 every group object in the $\infty$-category of ${\mathrm{Lie}}(1)$-algebras admits a unique delooping identifying group objects in ${\mathrm{Lie}}(1)$-algebras with $L_{\infty}$-algebras. This way we can identify the left hand side of adjunction (5) with $L_{\infty}$-algebras and obtain an adjunction
(6) ${\mathcal{U}}:\mathrm{Alg}_{\mathrm{Lie}}({\mathbb{Q}})\rightleftarrows{\mathrm{Hopf}}({\mathbb{Q}}):{\mathrm{Prim}}.$
@@
Inspired by the classical situation it seems reasonable to view the left adjoint ${\mathcal{U}}$ of the functor of derived primitive elements as a model for the functor of derived universal enveloping Hopf algebra. This model was identified by Gaitsgory-Rozenblyum with the classical definition [11, Theorem 6.1.2.] and fruitfully generalized by Knudsen [22] to stable homotopy theory to produce enveloping ${\mathbb{E}}_{\mathrm{n}}$-algebras replacing loops by n-fold loops.
@@
Motivated by these results we apply Koszul duality to construct and study a derived universal enveloping Hopf algebra for algebras over Ching’s spectral Lie operad [5] in any stable presentably symmetric monoidal $\infty$-category. Lurie [23, 5.2.] constructs a Koszul duality between augmented associative algebras and coaugmented coassociative coalgebras in any monoidal $\infty$-category that admits the necessary limits and colimits to perform the construction. Applied to the monoidal $\infty$-category of symmetric sequences Lurie’s Koszul duality gives rise to a Koszul duality between augmented non-unital $\infty$-operads and coaugmented non-counital $\infty$-cooperads [17, §4.3.]. Based on this operadic Koszul duality Amabel [1], Branter-Campos-Nuiten [3], Ching-Harper [4] and Francis-Gaitsgory [8] describe and study Koszul duality between algebras over any augmented non-unital $\infty$-operad ${\mathcal{O}}$ and coalgebras over the Koszul dual $\infty$-cooperad ${\mathcal{O}}^{\vee}$ in any stable presentably symmetric monoidal $\infty$-category ${\mathcal{C}}.$ This Koszul duality takes the form of an adjunction
(7) $\mathrm{TQ}_{\mathcal{O}}:\mathrm{Alg}_{\mathcal{O}}({\mathcal{C}})\rightleftarrows{\mathrm{coAlg}}^{\mathrm{dp},\hskip 0.56905pt{\mathrm{conil}}}_{{\mathcal{O}}^{\vee}}({\mathcal{C}})$
between the $\infty$-category of ${\mathcal{O}}$-algebras in ${\mathcal{C}}$ and the $\infty$-category of conilpotent ${\mathcal{O}}^{\vee}$-coalgebras that are equipped with divided powers. The left adjoint takes topological Quillen homology. Forgetting divided powers and conilpotence gives an adjunction
$\mathrm{TQ}_{\mathcal{O}}:\mathrm{Alg}_{\mathcal{O}}({\mathcal{C}})\rightleftarrows{\mathrm{coAlg}}_{{\mathcal{O}}^{\vee}}({\mathcal{C}}):{\mathrm{Prim}}_{\mathcal{O}}$
between the $\infty$-category of ${\mathcal{O}}$-algebras in ${\mathcal{C}}$ and the $\infty$-category of ${\mathcal{O}}^{\vee}$-coalgebras in ${\mathcal{C}}$, where the right adjoint takes a derived variant of primitive elements. Specializing to the Koszul duality [5] between the shifted spectral Lie operad ${\mathrm{Lie}}(1)$ and the non-unital cocommutative cooperad the latter adjunction specializes to an adjunction
(8) $\mathrm{TQ}_{{\mathrm{Lie}}(1)}:\mathrm{Alg}_{{\mathrm{Lie}}(1)}({\mathcal{C}})\rightleftarrows{\mathrm{coAlg}}_{{\mathbb{E}}_{\infty}}({\mathcal{C}}):{\mathrm{Prim}}_{{\mathrm{Lie}}(1)}$
between ${\mathrm{Lie}}(1)$-algebras and coaugmented ${\mathbb{E}}_{\infty}$-coalgebras generalizing adjunction (3) to stable homotopy theory. By Corollary 5.8 every group object in the $\infty$-category of ${\mathrm{Lie}}(1)$-algebras in ${\mathcal{C}}$ admits a unique delooping identifying group objects in ${\mathrm{Lie}}(1)$-algebras in ${\mathcal{C}}$ with $L_{\infty}$-algebras in ${\mathcal{C}}.$ So adjunction (8) induces on group objects an adjunction
@@
(9) ${\mathrm{Hopf}}({\mathcal{C}}):={\mathrm{Grp}}({\mathrm{coAlg}}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})):{\mathrm{Prim}}:={\mathrm{Prim}}_{{\mathrm{Lie}}(1)}.$
The latter adjunction extends adjunction (6) to stable homotopy theory and gives a model for the derived universal enveloping Hopf algebra.
@@
$\to\tau_{{\mathrm{n}}}({\mathcal{O}})\to\tau_{{\mathrm{n}}-1}({\mathcal{O}})\to...\to\tau_{1}({\mathcal{O}})={\mathrm{triv}}$
consisting of operadic truncations ${\mathcal{O}}\to\tau_{{\mathrm{n}}}({\mathcal{O}})$, where $\tau_{{\mathrm{n}}}({\mathcal{O}})_{\ell}:=\left\{\begin{array}[]{ll}{\mathcal{O}}_{\ell},&\ell\leq{\mathrm{n}}\\ 0,&\textrm{else}.\\ \end{array}\right.$ The latter tower gives for every ${\mathcal{O}}$-algebra ${\mathrm{X}}$ a tower of ${\mathcal{O}}$-algebras
$\to\tau_{{\mathrm{n}}}({\mathcal{O}})\circ_{\mathcal{O}}{\mathrm{X}}\to\tau_{{\mathrm{n}}-1}({\mathcal{O}})\circ_{\mathcal{O}}{\mathrm{X}}\to...\to{\mathrm{triv}}\circ_{\mathcal{O}}{\mathrm{X}}\simeq\mathrm{TQ}_{\mathcal{O}}({\mathrm{X}}),$
where $\circ$ is the composition product, interpolating between $\mathrm{TQ}_{\mathcal{O}}({\mathrm{X}})$ and the sequential limit $\lim_{{\mathrm{n}}\geq 1}\tau_{{\mathrm{n}}}({\mathcal{O}})\circ_{\mathcal{O}}{\mathrm{X}}.$ We prove the following:
Proposition 1.2.
(Proposition 4.13) Let ${\mathcal{C}}$ be a stable presentably symmetric monoidal $\infty$-category, ${\mathcal{O}}$ a non-unital $\infty$-operad in ${\mathcal{C}}$ such that ${\mathcal{O}}_{1}={\mathbb{1}}$ and ${\mathrm{X}}$ an ${\mathcal{O}}$-algebra in ${\mathcal{C}}$. There is a commutative square:
${{\mathrm{X}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${{\mathrm{Prim}}_{\mathcal{O}}{\mathrm{T}}{\mathrm{Q}}_{\mathcal{O}}({\mathrm{X}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\lim_{{\mathrm{n}}\geq 1}(\tau_{\mathrm{n}}({\mathcal{O}})\circ_{{\mathcal{O}}}{\mathrm{X}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\lim_{{\mathrm{n}}\geq 1}((\tau_{\mathrm{n}}({\mathcal{O}})\circ_{{\mathcal{O}}}{\mathrm{triv}})*^{{\mathcal{O}}^{\vee}}({\mathrm{triv}}\circ_{\mathcal{O}}{\mathrm{X}})),}$
We use Proposition 1.2 to offer an alternative proof of the following derived Milnor-Moore theorem that was proven by Gaitsgory-Rozenblyum [11, Theorem 4.4.6.] via a filtration of the primitive elements:
@@
Moreover we deduce the classical Milnor-Moore theorem (Theorem 5.36) from Theorem 1.3, which gives an alternative proof of the classical Milnor-Moore theorem.
Passing to abelian Hopf algebras Theorem 1.3 gives the following corollary:
Corollary 1.4.
(Corollary 5.32) Let ${\mathcal{C}}$ be a ${\mathbb{Q}}$-linear stable presentably symmetric monoidal $\infty$-category. The free ${\mathbb{E}}_{\infty}$-algebra functor
lifts to an embedding ${\mathcal{C}}\to{\mathrm{Hopf}}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})$ into the $\infty$-category of commutative and cocommutative Hopf algebras.
If ${\mathcal{C}}$ is not ${\mathbb{Q}}$-linear, Corollary 1.4 is not valid. In this case we offer the following remedy:
Proposition 1.5.
(Proposition 5.34) Let ${\mathcal{C}}$ be a stable presentably symmetric monoidal $\infty$-category. The free ${\mathbb{E}}_{\infty}$-algebra functor
lifts to an embedding ${\mathcal{C}}\to{\mathrm{Hopf}}^{\mathrm{dp},{\mathrm{conil}}}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})$ into the $\infty$-category of conilpotent cocommutative coalgebras equipped with divided powers.
By the Milnor-Moore theorem the unit of adjunction (1) is fully faithful. Therefore, also the map ${\mathcal{L}}\to{\mathcal{P}}{\mathcal{U}}{\mathcal{L}}\cong{\mathcal{P}}{\mathrm{T}}$ is an isomorphism, where ${\mathcal{L}}$ is the free Lie algebra and ${\mathrm{T}}={\mathcal{U}}{\mathcal{L}}$ is the tensor algebra. This implies that the map from the free Lie algebra monad ${\mathcal{L}}$ to the monad ${\mathcal{P}}\circ{\mathrm{T}}$ associated to the adjunction
${\mathrm{T}}={\mathcal{U}}\circ{\mathcal{L}}:\mathrm{Vect}_{K}\rightleftarrows{\mathrm{Lie}}_{K}\rightleftarrows{\mathrm{Hopf}}_{K}$
is an isomorphism. Consequently, by the Milnor-Moore theorem Lie algebras agree with algebras over the monad ${\mathcal{P}}\circ{\mathrm{T}}$ associated to the latter adjunction. Similarly, Theorem 1.3 implies the following corollary:
Corollary 1.6.
(Corollary 5.27) Let ${\mathcal{C}}$ be a ${\mathbb{Q}}$-linear stable presentably symmetric monoidal $\infty$-category. The $\infty$-category of Lie algebras in ${\mathcal{C}}$ is canonically equivalent to the $\infty$-category of algebras over the monad associated to the adjunction ${\mathrm{T}}={\mathcal{U}}\circ{\mathcal{L}}:{\mathcal{C}}\rightleftarrows{\mathrm{Hopf}}({\mathcal{C}}):{\mathrm{Prim}}.$
We extend the latter to a possible definition of a variant of Lie algebra: we construct for every additive presentably symmetric monoidal $\infty$-category ${\mathcal{C}}$ an adjunction
${\mathrm{T}}:{\mathcal{C}}\rightleftarrows{\mathrm{Hopf}}({\mathcal{C}}):{\mathrm{Prim}}$
between ${\mathcal{C}}$ and the $\infty$-category of cocommutative Hopf algebras in ${\mathcal{C}},$ where the left adjoint lifts the free associative algebra and the right adjoint assigns a derived version of primitive elements. We define a variant of Lie algebras, which we call restricted $L_{\infty}$-algebras, as algebras over the monad associated to the latter adjunction. For ${\mathcal{C}}$ the category of (differentially) (graded) ${\mathrm{K}}$-vector spaces this definition of restricted $L_{\infty}$-algebra recovers the notion of (differentially) (graded) restricted Lie algebra over ${\mathrm{K}}$ by a result of Fresse [10, Theorem 1.2.5.] or likewise a version of the Milnor-Moore theorem in positive characteristic.
We prove the following theorem:

### v2

# A derived Milnor-Moore theorem

@@ Abstract.
For every stable presentably symmetric monoidal $\infty$-category $\mathcal{C}$} we use the Koszul duality between the spectral Lie operad and the cocommutative cooperad to construct an enveloping Hopf algebra functor ${\mathcal{U}}:\mathrm{Alg}_{\mathrm{Lie}}({\mathcal{C}})\to{\mathrm{Hopf}}({\mathcal{C}})$ from Lie algebras in $\mathcal{C}$} to cocommutative Hopf algebras in $\mathcal{C}$} left adjoint to a functor of derived primitive elements $\mathrm{Prim}$}. We study the unit of this adjunction in rational and chromatic homotopy theory: we prove that if $\mathcal{C}$} is a rational stable presentably symmetric monoidal $\infty$-category, the enveloping Hopf algebra functor ${\mathcal{U}}:\mathrm{Alg}_{\mathrm{Lie}}({\mathcal{C}})\to{\mathrm{Hopf}}({\mathcal{C}})$ is fully faithful reproving a result of Gaitsgory-Rozenblyum. Let $n\geq 1$ be a natural and $\Phi[-1]:{\mathcal{S}}_{v_{n}}\to\mathrm{Alg}_{\mathrm{Lie}}({\mathrm{Sp}}_{T_{n}})$ the shifted Bousfield-Kuhn functor from $v_{n}$-periodic homotopy types to spectral Lie algebras in $T_{n}$-local spectra. We prove that for every $v_{n}$-periodic homotopy type $X$ the unit $\Phi(X)[-1]\to{\mathrm{Prim}}{\mathcal{U}}(\Phi(X)[-1])$ identifies with the Goodwillie completion $\Phi\to\lim_{n\geq 0}P_{n}(\Phi)$ evaluated at the loop space of $X.$
@@ 1. Introduction
In the seminal paper [40] Milnor and Moore prove that every graded Lie algebra over a field of characteristic zero arises as the primitive elements of its universal enveloping Hopf algebra [40, Proposition 5.17.]. From this result Milnor and Moore reveal a fundamental connection between Lie algebras and Hopf algebras over any field $K$ of characteristic zero. Forming the universal enveloping Hopf algebra and forming the primitive elements are part of an adjunction
(1.1) ${\mathcal{U}}:{\mathrm{Lie}}_{K}\rightleftarrows{\mathrm{Hopf}}_{K}:{\mathrm{Prim}}$
between the categories of graded Lie algebras and graded Hopf algebras over $K$. By the Milnor-Moore theorem the unit of this adjunction is invertible, which implies that the enveloping Hopf algebra functor $\mathcal{U}$} is fully faithful. Moreover Milnor-Moore [40, Theorem 5.18.] characterize the Hopf algebras in the essential image of the enveloping Hopf algebra $\mathcal{U}$} as the primitively generated Hopf algebras, those graded Hopf algebras whose underlying graded algebra is generated by the subspace of primitive elements.
@@
(1.2) ${\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces{\mathrm{Ho}}({\mathcal{S}}^{\geq 2}_{{\mathbb{Q}}})}$$\scriptstyle{{\mathcal{C}}}\scriptstyle{{\mathcal{L}}}$${{\mathrm{Ho}}(\mathrm{co}\mathrm{Alg}({\mathbb{Q}})^{\geq 2})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${{\mathrm{Ho}}({\mathrm{Lie}}({\mathbb{Q}})^{\geq 1})}$
@@
Quillen builds the Lie model of a simply connected space $X$ in three steps. First he forms Kan’s loop group, a simplicial group that endows the loop space $\Omega({\mathrm{X}})$ with a group structure. As a second step he applies rational chains to Kan’s loop group to obtain a simplicial rational cocommutative Hopf algebra. Finally, he takes the primitive elements of this simplicial Hopf algebra to obtain a simplicial rational Lie algebra, which corresponds to a rational differentially graded Lie algebra via Dold-Kan’s correspondence. Quillen’s construction guarantees that the homology of the Lie model identifies with the primitive elements of the rational homology of the loop space, in symbols $H_{*}({\mathcal{L}}(X))\cong{\mathrm{Prim}}(H_{*}(\Omega(X);{\mathbb{Q}})).$ Since $X$ is simply connected, $H_{*}(\Omega(X);{\mathbb{Q}})$ is connected, which ensures that the graded Hopf algebra $H_{*}(\Omega(X);{\mathbb{Q}})$ is primitively generated. As a consequence the Milnor-Moore theorem identifies the enveloping Hopf algebra of
$\pi_{*}(\Omega(X))\otimes{\mathbb{Q}}\cong H_{*}({\mathcal{L}}(X))\cong{\mathrm{Prim}}(H_{*}(\Omega(X);{\mathbb{Q}}))$
with $H_{*}(\Omega(X);{\mathbb{Q}})$ and so reveals a fundamental relationship between rational homotopy and rational homology of a loop space.
[…]
(1.3) $\mathrm{Alg}_{{\mathrm{Lie}}(1)}({\mathbb{Q}})\rightleftarrows\mathrm{co}\mathrm{Alg}_{{\mathbb{E}}_{\infty}}({\mathbb{Q}}):{\mathrm{Prim}}$
[…]
(1.4) $\mathrm{Alg}_{{\mathrm{Lie}}}({\mathbb{Q}})^{\geq 1}\simeq\mathrm{co}\mathrm{Alg}_{{\mathbb{E}}_{\infty}}({\mathbb{Q}})^{\geq 2}$
[…]
(1.5) ${\mathrm{Grp}}(\mathrm{Alg}_{{\mathrm{Lie}}(1)}({\mathbb{Q}}))\rightleftarrows{\mathrm{Hopf}}({\mathbb{Q}}):={\mathrm{Grp}}(\mathrm{co}\mathrm{Alg}_{{\mathbb{E}}_{\infty}}({\mathbb{Q}})):{\mathrm{Prim}}.$
@@
By Corollary 4.8 every group object in the $\infty$-category of ${\mathrm{Lie}}(1)$-algebras admits a unique delooping identifying group objects in ${\mathrm{Lie}}(1)$-algebras with $L_{\infty}$-algebras. This way we can identify the left hand side of adjunction (1.5) with $L_{\infty}$-algebras and obtain an adjunction
(1.6) ${\mathcal{U}}:\mathrm{Alg}_{\mathrm{Lie}}({\mathbb{Q}})\rightleftarrows{\mathrm{Hopf}}({\mathbb{Q}}):{\mathrm{Prim}}.$
@@
Inspired by the classical situation it seems reasonable to view the left adjoint $\mathcal{U}$} of the functor of derived primitive elements as a model for the functor of derived universal enveloping Hopf algebra. This model was identified by Gaitsgory-Rozenblyum with the classical definition [18, Theorem 6.1.2.] and generalized by Brantner-Camarena-Heuts [9], Knudsen [35] and Shi [48] to stable homotopy theory to produce enveloping ${\mathbb{E}}_{n}$-algebras replacing loops by n-fold loops.
@@
Motivated by these results we apply Koszul duality to construct and study a derived universal enveloping Hopf algebra for algebras over Ching’s spectral Lie operad [11] in any stable presentably symmetric monoidal $\infty$-category. Lurie [39, 5.2.] constructs Koszul duality between augmented associative algebras and coaugmented coassociative coalgebras in any monoidal $\infty$-category that admits the necessary limits and colimits to perform the construction. Applied to the monoidal $\infty$-category of symmetric sequences Lurie’s Koszul duality gives rise to a Koszul duality between augmented non-unital $\infty$-operads and coaugmented non-counital $\infty$-cooperads [22, §4.3.]. Based on this operadic Koszul duality Amabel [1], Branter-Campos-Nuiten [8], Ching-Harper [10] and Francis-Gaitsgory [15] describe and study Koszul duality between algebras over any augmented non-unital $\infty$-operad $\mathcal{O}$} and coalgebras over the Koszul dual $\infty$-cooperad $\mathcal{O}$$^{\vee}$ in any stable presentably symmetric monoidal $\infty$-category $\mathcal{C}$}. This Koszul duality takes the form of an adjunction
(1.7) $\mathrm{TQ}_{\mathcal{O}}:\mathrm{Alg}_{\mathcal{O}}({\mathcal{C}})\rightleftarrows\mathrm{co}\mathrm{Alg}^{\mathrm{dp},\hskip 0.56905pt{\mathrm{conil}}}_{{\mathcal{O}}^{\vee}}({\mathcal{C}})$
[…]
$\mathrm{TQ}_{\mathcal{O}}:\mathrm{Alg}_{\mathcal{O}}({\mathcal{C}})\rightleftarrows\mathrm{co}\mathrm{Alg}_{{\mathcal{O}}^{\vee}}({\mathcal{C}}):{\mathrm{Prim}}_{\mathcal{O}}$
[…]
(1.8) $\mathrm{TQ}_{{\mathrm{Lie}}(1)}:\mathrm{Alg}_{{\mathrm{Lie}}(1)}({\mathcal{C}})\rightleftarrows\mathrm{co}\mathrm{Alg}_{{\mathbb{E}}_{\infty}}({\mathcal{C}}):{\mathrm{Prim}}_{{\mathrm{Lie}}(1)}$
[…]
@@
(1.9) ${\mathrm{Hopf}}({\mathcal{C}}):={\mathrm{Grp}}(\mathrm{co}\mathrm{Alg}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})):{\mathrm{Prim}}:={\mathrm{Prim}}_{{\mathrm{Lie}}(1)}.$
The latter adjunction extends adjunction (1.6) to stable homotopy theory and gives a model for the derived universal enveloping Hopf algebra.
@@
$\to\tau_{{n}}({\mathcal{O}})\to\tau_{{n}-1}({\mathcal{O}})\to...\to\tau_{1}({\mathcal{O}})={\mathrm{triv}}$
[…]
$\to\tau_{{n}}({\mathcal{O}})\circ_{\mathcal{O}}{\mathrm{X}}\to\tau_{{n}-1}({\mathcal{O}})\circ_{\mathcal{O}}{\mathrm{X}}\to...\to{\mathrm{triv}}\circ_{\mathcal{O}}{\mathrm{X}}\simeq\mathrm{TQ}_{\mathcal{O}}({\mathrm{X}}),$
where $\circ$ is the composition product, interpolating between $\mathrm{TQ}_{\mathcal{O}}({\mathrm{X}})$ and the $\mathrm{TQ}_{\mathcal{O}}$-completion $\mathrm{TQ}_{\mathcal{O}}(X)^{\wedge}:=\lim_{{n}\geq 1}\tau_{{n}}({\mathcal{O}})\circ_{\mathcal{O}}{\mathrm{X}}.$ We prove the following theorem:
Theorem 1.1.
(Theorem 4.17) Let $\mathcal{C}$} be a stable presentably symmetric monoidal $\infty$-category such that all norm maps of objects in $\mathcal{C}$} with an action of a symmetric group are equivalences. Let $\mathrm{X}$} be a Lie algebra in $\mathcal{C}$}. The unit $X\to{\mathrm{Prim}}{\mathcal{U}}(X)$ identifies with the completion map
$X\to\mathrm{TQ}_{\mathrm{Lie}}(X)^{\wedge}.$
Before we explain how we prove Theorem˜1.1, we explain important corollaries. We apply Theorem 1.1 to rational and chromatic homotopy theory. Rationally, we immediately obtain the following derived Milnor-Moore theorem that was proven by Gaitsgory-Rozenblyum [18, Theorem 4.4.6.] via an elaborate filtration of the primitive elements.
@@
Moreover we explain how to deduce the classical Milnor-Moore theorem (Theorem 4.39) from Theorem 1.2, which gives an alternative proof of the classical version.
We apply Theorem 1.1 to chromatic homotopy theory. Chromatic homotopy theory [13], [33], [45], [46], [47] is an extension of rational homotopy. For a fixed prime $p$ and height $n\geq 1$ one considers $v_{n}$-periodic homotopy groups, an extension of rational homotopy groups that captures periodic phenomena in homotopy groups. Extending the concept of rational equivalences chromatic homotopy theory studies $v_{n}$-periodic equivalences, those maps of pointed $p$-local spaces inducing an equivalence on $v_{n}$-periodic homotopy groups. Extending the concept of rational spaces chromatic homotopy theory studies $p$-local pointed spaces up to $v_{n}$-periodic equivalences, i.e. the Bousfield localization ${\mathcal{S}}_{(p)}\rightleftarrows{\mathcal{S}}_{v_{n}}$ of the category ${\mathcal{S}}_{(p)}$ of pointed $p$-local spaces with respect to the class of $v_{n}$-periodic equivalences. The latter exists by [30, Theorem 2.2.]. Similarly, one considers the stable homotopy theory of $v_{n}$-periodic homotopy types, which by [30, Remark 3.20.] is also known as the homotopy theory of $T_{n}$-local spectra . The $v_{n}$-periodic homotopy groups are elegantly described via the Bousfield-Kuhn functor [5], [6], [7], which is a key player in chromatic homotopy theory [36], [3], [30].
The latter is a functor $\Phi:{\mathcal{S}}_{(p)}\to{\mathrm{Sp}}_{T_{n}}$ from $p$-local pointed spaces to $T_{n}$-local spectra that assigns to a $p$-local space $X$ a $T_{n}$-local spectrum $\Phi(X)$ whose homotopy groups are the $v_{n}$-periodic homotopy groups of $X$, and such that for every $p$-local spectrum $Y$ the $T_{n}$-local spectrum $\Phi(\Omega^{\infty}(Y))$ is the $T_{n}$-localization of $Y$ (see [49, Part I] for details).
Heuts [30, Theorem 2.6.] extends Quillen’s rational homotopy theory [44] to higher chromatic heights. Heuts proves that every $v_{n}$-periodic homotopy type $X$ is determined by a shifted Lie structure on the $T_{n}$-local spectrum $\Phi(X)$ in analogy that every simply connected rational space is determined by a rational differently graded Lie algebra. Precisely, Heuts proves that the Bousfield-Kuhn functor $\Phi:{\mathcal{S}}_{v_{n}}\to{\mathrm{Sp}}_{T_{n}}$ lifts to an equivalence ${\mathcal{S}}_{v_{n}}\to\mathrm{Alg}_{{\mathrm{Lie}}(1)}({\mathrm{Sp}}_{T_{n}})$. Composing with the equivalence $\mathrm{Alg}_{{\mathrm{Lie}}(1)}({\mathrm{Sp}}_{T_{n}})\simeq\mathrm{Alg}_{{\mathrm{Lie}}}({\mathrm{Sp}}_{T_{n}})$ that loops the underlying spectrum, one obtains an equivalence $\Phi[-1]:{\mathcal{S}}_{v_{n}}\to\mathrm{Alg}_{{\mathrm{Lie}}}({\mathrm{Sp}}_{T_{n}})$.
We deduce the following chromatic variant of the Milnor-Moore theorem: